calculate the pH of a 1.3M weak acid solution that is 2.8% ionized
Answer #1

Molarity of weak Acid (HA) = 1.3M

HA ------------ H+ + A-

Acid is ionized 2.8% i.e.

Molarity of H+ ion = [H+] = (1.3*2.8)/100 = 0.0364M

pH = -log[H+] = -log(0.0364) = 1.439

pH = 1.439 ...... Answer